# What Lies Beneath? Exploring the Impact of Underlying AI Model Updates in AI-Infused Systems

### v2

@@ Abstract.
AI models are constantly evolving, with new versions released frequently. This raises a key question: how should AI-infused systems integrate updates when the downstream impact on user experience and performance is unclear? Human-AI interaction guidelines encourage notifying users about (changes in) model capabilities, ideally supported by thorough benchmarking. Yet, as AI models integrate into domain-specific workflows, exhaustive benchmarking can become impractical or expensive, often leading to invisible or minimally communicated updates. In this work, we explore the impact of such updates through two complementary studies on facial recognition for historical person identification. First, we conducted an online experiment to understand how users distinguish between models, followed by a diary study examining user perceptions in a real-world deployment. Our findings reveal how model changes impact human-AI performance, downstream user behavior, and the folk theories they develop. Based on these insights, we discuss implications for updating models in AI-infused systems.
AI model update, facial recognition, user behavior, user perception of AI models, information retrieval, transparency, human-AI interaction guidelines, historical person identification
@@ 1. Introduction
AI model developers are pushing out new models and updates at a rapid pace, driven by growing user needs and the necessity to incorporate newer capabilities and address inaccuracies. For instance, Hugging Face hosts over 300,000 models (Tousignant, 2023), and in 2023 alone, 149 new foundation models were released (Maslej et al., 2024). These models power a wide range of AI-infused applications across domains such as healthcare, finance, social media, and autonomous vehicles (Maslej et al., 2024). As these underlying models get updated, a critical question arises: how should the applications that build upon them incorporate these changes, and how do these updates impact the downstream user experience?
@@
Previous work has shown that model updates do not always improve human-AI team performance (Bansal et al., 2019b; Chattopadhyay et al., 2017) and can lead to user dissatisfaction due to misaligned expectations or disruptions in user workflows (Kocielnik et al., 2019; Bucher, 2012). Such unintended consequences can manifest in various ways, be it unforeseen accidents following updates to autonomous vehicle algorithms (Cao, 2023), or distress among social media users due to alterations in recommendation algorithms (Register et al., 2023), or diminished volunteer engagement in citizen science projects upon the integration of automated systems (Trouille et al., 2019). To avoid unexpected shifts in expectations, human-AI interaction guidelines advocate for notifying users about model changes and capabilities (Amershi et al., 2019).
@@
Effectively communicating an AI model’s capabilities requires a deep understanding of what the model can do, which in turn demands comprehensive benchmarking and stress testing across diverse scenarios. However, as these models are deployed in workflows or domains that developers did not originally anticipate or design for, the complexity of thorough benchmarking increases (Raji et al., 2021). This makes comprehensive testing both impractical and costly, leading to silent updates or minimal communication or approximation about changes and capabilities, which leaves users unaware of significant model modifications. When users are unaware of these changes, they may interact with the system based on outdated assumptions, leading to potential errors, inefficiencies, or frustrations (Bansal et al., 2019a; Villareale et al., 2022).
With the ubiquity of AI-infused applications, many lay end-users may not fully understand AI model capabilities, let alone grasp what has changed following updates. This lack of awareness raises key questions about how well users can adapt to these changes and whether they can complete their tasks effectively. In this work, we aim to understand whether users of an AI-infused application can detect changes in the underlying models in such practical scenarios with limited communication, and how these updates affect their performance, perceptions, and behavior.
@@
To explore these questions, we conducted two complementary studies involving a facial recognition-based system for completing historical person identification tasks (i.e., finding potential matches from a ranked list of search results for a given query photo): 1) a controlled experiment where users were unaware of the underlying face recognition model being updated, and 2) a real-world deployment where they could toggle between the old and new models.
@@
For the first study, we ran an online experiment to determine whether users could distinguish changes between facial recognition models across successive trials and what cues they relied on to do so. We also explored how the underlying model impacted their task performance, perceptions of accuracy, and behavior, particularly when they were unaware of the model change. We complemented this with a second study in a real-world setting on Civil War Photo Sleuth (CWPS) (Mohanty et al., 2019b), a widely-used platform for identifying historical photos through a facial recognition-based workflow. In April 2023, CWPS added a new facial recognition model alongside the existing one, allowing users to toggle between the two models. Through a two-week long diary study with 10 active users, we aimed to understand how they perceived the new model in contrast to the old one, and what factors influenced their model preferences.
@@
Our findings provide key insights into users’ abilities to detect model changes, the cues they rely on, and how these changes influence their task performance, perceptions of accuracy, behavior, and the folk theories and preferences they develop. We discuss the complexities involved in end-user evaluation of AI models and the implications for deploying model updates in a human-AI collaborative environment.
@@ 2.1. User Frustration with Software Updates
This resistance to updates is reflected in the findings from Rula et al. (Rula et al., 2020), which showed that many users continue to run outdated software, avoiding updates due to concerns about potential disruptions or changes. This behavior further demonstrates how user apprehension can hinder the adoption of updates, leading to increased security vulnerabilities and reduced system performance. While these studies highlight software updates, they parallel our focus on AI model updates, which further add unpredictability, especially with generative AI, making it critical to understand downstream effects with varying levels of user communication.
@@ 2.2. User Perceptions of Dynamic AI Systems
Advances in AI research have opened new avenues for human-AI collaboration, leveraging the complementary strengths of humans and AI systems to drive progress across diverse domains such as creativity, manufacturing, and journalism (Wilson and Daugherty, 2018; Liu et al., 2023; Broussard et al., 2019). However, this collaboration is far from static or fully understood. Both human users and AI agents operate dynamically, with interactions shaped by factors such as context of use, task complexity, and the continually evolving—yet often uncertain—nature of AI technologies (Yang et al., 2020; Amershi et al., 2019; Horvitz, 1999). These complexities are further amplified by the growing push toward Agentic AI, where large language model (LLM) agents are increasingly designed to act autonomously and make independent decisions (Wu et al., 2023; Maslej et al., 2024).
While the fusion of human expertise with AI capabilities has led to significant advancements, it is crucial to recognize the inherent complexities and potential drawbacks that accompany this collaboration. Dynamic human-AI interactions can give rise to challenges such as mistrust, misinformation, and data privacy concerns (Zhou et al., 2023; Spitale et al., 2023). In AI-assisted decision-making (Lai et al., 2023), additional pitfalls include automation bias, amplification of incorrect results, and the perpetuation of existing biases, which can further hinder harmonious collaboration (Willever-Farr and Forte, 2014; Jones-Jang and Park, 2023; Kim et al., 2024).
@@
As the market experiences a rapid influx of new AI models, this frequent turnover, while often beneficial, presents unique challenges. One of the key questions is how users perceive and adapt to changes in AI systems, particularly when updates to the underlying models are not clearly communicated. Prior work has explored how users may perceive shifts in model explanations after updates, even when the decisions remain unchanged (Wang and Yin, 2023). These shifts can affect users’ trust, particularly if the new explanations diverge significantly from previous ones. Users may experience cognitive dissonance when their established mental models are violated, leading to reduced trust and satisfaction with the system.
@@
In this work, we explore how users perceive AI model updates both in a controlled environment and a real-world deployment, focusing on factors such as response speed, the number of results generated, and the cognitive load required to interpret changes. By examining these elements under different levels of communication about model updates, we aim to provide practical validation for existing human-AI interaction guidelines (Amershi et al., 2019),
@@ 3.1. Hypotheses
Based on prior work and the distinctive differences between the two models (see Section 3.4), we propose the following hypotheses to guide our investigation:
  1. H1
    Users will be able to distinguish a change in the model even though they are interacting with black-box systems.
    Prior work has shown that users can detect subtle shifts in system behavior, though this has primarily been tested with model explanations in non-black-box, controlled settings (Wang and Yin, 2023).
  2. H2
    Users will rely on observable model characteristics such as the number of search results and latency to distinguish between models.
    Research shows that users often focus on tangible factors like speed and the volume of results while interacting with AI models (Park et al., 2019; Oulasvirta et al., 2009).
  3. H3
    Users will perceive the more accurate model (newer) as more accurate, even if they are not explicitly informed about model changes.
    Drawing from different cognitive science theories, users are likely to perceive the model that retrieves more similar-looking candidates as more accurate because they can detect subtle improvements in output quality by focusing on high-diagnostic and directly comparable features (Markman and Gentner, 1996; Tversky, 1977).
  4. H4
    Human-AI team performance will improve with the more accurate model.
    Higher model accuracy contributes to reduced errors, which in turn improves overall team outcomes (Towler et al., 2023). However, prior work has also shown that may not always be true (Bansal et al., 2019b; Chattopadhyay et al., 2017).
  5. H5
    The underlying model will affect user behavior in terms of how many results they check and how they make decisions.
    Changes in recommender algorithms, as seen in social media platforms, have been shown to significantly alter user behavior, leading them to engage more or less with certain types of content (Owen, 2019; Karizat et al., 2021). Users adjust their interaction patterns based on perceived changes in output quality, relevance, or volume, influencing their decision-making and exploration of results.
  6. H6
    When users believe that a model has switched, it will lead to changes in their behavior.
@@
    According to Cognitive Dissonance Theory (Festinger, 1957), when users experience a mismatch between their expectations and system behavior, they adjust their actions to reduce the discomfort caused by this inconsistency, leading to changes in their decision-making and interaction strategies.
@@ 3.2. Experiment Setup
To address these questions, we designed an online survey experiment where participants interacted with a facial recognition-based web application (see Figure 1), modeled after the search results page on CWPS (Mohanty et al., 2019b). In each trial, participants were shown a query photo and tasked with identifying potential matches (i.e., other photos of the same person) from a ranked list of search results retrieved by a black-box facial recognition model. Participants visually inspected the results to determine whether any of the retrieved images matched the person in the query photo.
@@
Each participant completed 8 trials, with each trial featuring a different query photo and its corresponding search results (i.e., potential candidates). We randomized the underlying model used in each trial (i.e., the model could either remain the same or change between trials). Participants were not informed of any model changes, allowing us to investigate whether they could detect changes and what cues they relied on. This design enabled us to examine shifts in user perceptions, performance, and behaviors across trials.
3.3. Experiment Interface
We designed the web application to guide participants through the person identification tasks across eight consecutive trials. The interface supported five key components:

### v5

@@ Abstract.
AI models are constantly evolving, with new versions released frequently. Human-AI interaction guidelines encourage notifying users about changes in model capabilities, ideally supported by thorough benchmarking. However, as AI systems integrate into domain-specific workflows, exhaustive benchmarking can become impractical, often resulting in silent or minimally communicated updates. This raises critical questions: Can users notice these updates? What cues do they rely on to distinguish between models? How do such changes affect their behavior and task performance? We address these questions through two studies in the context of facial recognition for historical photo identification: an online experiment examining users’ ability to detect model updates, followed by a diary study exploring perceptions in a real-world deployment. Our findings highlight challenges in noticing AI model updates, their impact on downstream user behavior and performance, and how they lead users to develop divergent folk theories. Drawing on these insights, we discuss strategies for effectively communicating model updates in AI-infused systems.
AI Model Updates, Crowdsourcing, Facial Recognition, User Perception of AI Models, User Behavior, Historical Photo Identification, Folk Theories, Diary Study, Quantitative Methods
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713751isbn: 979-8-4007-1394-1/25/04ccs: Information systems Users and interactive retrievalccs: Computing methodologies Computer vision tasksccs: Human-centered computing Laboratory experimentsccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
AI model developers are pushing out new models and updates at a rapid pace, driven by growing user needs and the necessity to incorporate newer capabilities and address inaccuracies. For instance, Hugging Face hosts over 300,000 models (Tousignant, 2023), and in 2023 alone, 149 new foundation models were released (Maslej et al., 2024). These models power a wide range of AI-infused applications across domains such as healthcare, finance, social media, and autonomous vehicles (Maslej et al., 2024). However, with the ubiquity of AI-infused applications, many lay end-users may not fully understand AI model capabilities (Kennedy et al., 2023; Nader et al., 2024), let alone recognize or comprehend changes introduced by updates. This lack of awareness raises important questions: Can users notice updates in AI models? What cues do they rely on to infer changes in underlying AI models? How do such updates influence their behavior, perceptions, and task performance?
@@
Understanding how users perceive and adapt to these updates is critical, as updates often aim to improve performance but can unintentionally disrupt workflows, misalign expectations, or even lead to significant user dissatisfaction. Prior work has shown that model updates do not always improve human-AI team performance (Bansal et al., 2019b; Chattopadhyay et al., 2017) and can lead to user dissatisfaction due to misaligned expectations or disruptions in user workflows (Kocielnik et al., 2019; Bucher, 2012). Such unintended consequences can manifest in various ways, be it unforeseen accidents following updates to autonomous vehicle algorithms (Cao, 2023), or distress among social media users due to alterations in recommendation algorithms (Register et al., 2023), or diminished volunteer engagement in citizen science projects upon the integration of automated systems (Trouille et al., 2019). To avoid unexpected shifts in expectations, human-AI interaction guidelines advocate for notifying users about model changes and capabilities (Amershi et al., 2019).
@@
Effectively communicating an AI model’s capabilities requires a deep understanding of what the model can and cannot do, which in turn demands comprehensive benchmarking and stress testing across diverse scenarios. However, as these models are deployed in workflows or domains that developers did not originally anticipate or design for, the complexity of thorough benchmarking increases (Raji et al., 2021). This makes comprehensive testing both impractical and costly, leading to silent updates or minimal communication or approximation about changes and capabilities, which leaves users unaware of significant model modifications. When users are unaware of these changes, they may interact with the system based on outdated assumptions, leading to potential errors, inefficiencies, or frustrations (Bansal et al., 2019a; Villareale et al., 2022).
@@
To understand how users perceive and adapt to AI model updates in practical scenarios, we studied a facial recognition-based system used for historical person identification tasks (i.e., finding potential matches from a ranked list of search results for a given query photo). We conducted two complementary studies: 1) a controlled experiment simulating user interactions without explicit awareness of model updates, and 2) a real-world deployment where users could toggle between the old and new models.
@@
For the first study, we ran an online experiment on Prolific to determine whether users could distinguish changes between facial recognition models across successive trials and what cues they relied on to do so. We also explored how the underlying model impacted their task performance, perceptions of accuracy and behavior, particularly when they were unaware of the model change. We complemented this with a second study in a real-world setting on Civil War Photo Sleuth (CWPS) (Mohanty et al., 2019b), a widely-used platform for identifying historical photos through a facial recognition-based workflow. In April 2023, CWPS added a new facial recognition model alongside the existing one, allowing users to toggle between the two models. Through a two-week long diary study with 10 active users, we aimed to understand how they perceived the new model in contrast to the old one, and what factors influenced their model preferences.
@@
Our findings provide key insights into users’ abilities to detect model changes, the cues they rely on, and how these changes influence their task performance, perceptions of accuracy, behavior, and the folk theories and preferences they develop. Based on these findings, we discuss implications for deploying model updates in a human-AI collaborative environment.
@@ 2.1. User Frustration with Software Updates
This resistance to updates is reflected in the findings from Rula et al. (Rula et al., 2020), which showed that many users continue to run outdated software, avoiding updates due to concerns about potential disruptions or changes. This behavior further demonstrates how user apprehension can hinder the adoption of updates, leading to increased security vulnerabilities and reduced system performance. While these studies highlight software updates, AI model updates introduce new challenges that differ significantly from traditional software changes. Unlike software updates, which often involve visible changes to functionality or interfaces, AI model updates can occur silently, with changes in behavior or performance that are subtle or opaque to users. These updates, especially in black-box systems, may not provide clear cues that allow users to recognize when a change has occurred.
In our studies, we explore whether users can notice changes in the underlying AI model, and if so, what cues they rely on, and how they characterize the differences between the models. Understanding whether users can distinguish such updates is critical, as the inability to do so might lead to similar downstream effects on user trust, performance, or decision-making.
@@ 2.2. User Perceptions of Dynamic AI Systems
Advances in AI research have opened new avenues for human-AI collaboration, leveraging the complementary strengths of humans and AI systems to drive progress across diverse domains such as creativity, manufacturing, and journalism (Wilson and Daugherty, 2018; Liu et al., 2023; Broussard et al., 2019). However, this collaboration is far from static or fully understood. Human users and AI agents do not interact in fixed ways; instead, their behaviors and interactions evolve dynamically, influenced by factors such as context of use, task complexity, system outputs, and the continually evolving, yet often uncertain, nature of AI technologies (Yang et al., 2020; Amershi et al., 2019; Horvitz, 1999; Shavit et al., 2023).
AI-assisted decision-making — an example of human-AI teaming, where humans make decisions based on AI-provided suggestions and recommendations (Bansal et al., 2019a; Berretta et al., 2023; Lai et al., 2023) — illustrates this interplay. While such collaboration has led to significant advancements, it also introduces challenges. In these workflows, pitfalls such as automation bias, amplification of incorrect results, and the perpetuation of existing biases can undermine trust and effectiveness, highlighting the need for a more nuanced understanding of human-AI dynamics (Willever-Farr and Forte, 2014; Jones-Jang and Park, 2023; Kim et al., 2024).
@@
As the market experiences a rapid influx of new AI models (CNBC, 2024; Wiggers, 2024; Nuñez, 2024; Reuters, 2024b, a), this frequent turnover, while often beneficial, presents unique challenges. One of the key questions is how users perceive and adapt to changes in AI systems, particularly when updates to the underlying models are not clearly communicated. Prior work has explored how users may perceive shifts in model explanations after updates, even when the decisions remain unchanged (Wang and Yin, 2023). These shifts can affect users’ trust, particularly if the new explanations diverge significantly from previous ones. Users may experience cognitive dissonance when their established mental models are violated, leading to reduced trust and satisfaction with the system.
@@
In this work, we explore how users perceive AI model updates both in a controlled environment and a real-world deployment, focusing on the cues they rely on and how these updates influence their performance and behavior. Examining these elements in practical contexts of AI model updates can help in validating human-AI interaction guidelines, particularly those related to effectively communicating model changes (Amershi et al., 2019).
@@ 3.1. Hypotheses
We build on prior work and the key differences between the two models (see Section 3.3) to propose hypotheses guiding our investigation into whether users can detect changes in an underlying AI model, the cues they rely on, and how these updates influence their behavior and performance. Prior work has shown that users can detect subtle shifts in system behavior, primarily when provided with model explanations in non-black-box, controlled settings (Wang and Yin, 2023). Real-world cases, such as user discussions on Reddit about perceived changes in GitHub Copilot’s code suggestions and backlash to Replika updates that altered conversational behaviors (Reddit user, [n. d.]; Cole, 2023), indicate that users may notice changes in AI behavior even without explicit communication. Extending these findings, we hypothesize H1: Users will be able to distinguish a change in the model even though they are interacting with black-box systems.
We investigate these questions within the context of a historical person identification task where users examine AI-retrieved search results (i.e., a ranked list of visually similar candidates) and combine these with their own judgment to identify potential matches. In such retrieval tasks, like web search engines, observable cues like latency and the number of results displayed shape how users perceive and interact with the system, potentially serving as key indicators for distinguishing between models. Prior work demonstrates that users are sensitive to latency, with even small delays being noticed and impacting satisfaction (Brutlag, 2009; Teevan et al., 2013). Similarly, the number of results shapes user perception: too many options can overwhelm users, while fewer results increase focus but risk neglecting potentially relevant items (Oulasvirta et al., 2009; Kelly and Azzopardi, 2015). Building on these insights, we hypothesize: H2: Users will rely on observable system characteristics, such as the number of search results and latency, to distinguish between models.
In addition to examining cues for distinguishing between models, we also investigate how users perceive the accuracy of the underlying face recognition models. Based on cognitive science theories of similarity, users are likely to perceive the model that retrieves candidates that are more similar-looking as more accurate because they can detect subtle improvements in output quality by focusing on high-diagnostic and directly comparable features (Markman and Gentner, 1996; Tversky, 1977). Similarly, prior research in search engines suggests that fewer, more precise results — a trait of the newer face recognition model we tested (see Section 3.3) — can increase user trust in the system by reducing cognitive load and emphasizing relevance (Oulasvirta et al., 2009). These findings inform H3: When comparing two models (older vs. newer), users will perceive the newer model (deemed more accurate by developers) as more accurate, even if they are not explicitly aware of which model they are using or informed about model changes.
Updating an AI model to improve its standalone accuracy raises the question of whether these improvements translate to better human-AI team performance. While higher model accuracy intuitively reduces errors and enhances outcomes (Towler et al., 2023), prior research highlights that this may not always hold true, as mismatches between AI improvements and human workflows can hinder collaboration (Bansal et al., 2019b; Chattopadhyay et al., 2017). Despite these complexities, we hypothesize H4: Human-AI team performance will improve with the more accurate model.
Model updates in AI systems can introduce changes that subtly alter user interactions, prompting adjustments in behavior. Prior research on software updates shows that unexpected changes to functionality or outputs can lead users to modify their workflows, engagement levels, or decision-making processes (Morreale and Eriksson, 2020; Mathur et al., 2018). Similarly, changes in recommender algorithms, as observed in social media platforms, have been shown to influence user behavior by shifting how they explore or engage with content based on perceived output quality, relevance, or volume (Owen, 2019; Karizat et al., 2021). In the context of a retrieval task, such shifts might manifest as differences in the number of results users check, the time spent on tasks, or their decision-making strategies. Building on these insights, we hypothesize: H5: AI model updates will affect user behavior in terms of time spent on tasks, the number of results they check, and the decisions they make.
@@
Moreover, when users believe a model has been updated, even if this belief is not explicitly confirmed, it can lead to behavioral changes driven by cognitive dissonance. According to Cognitive Dissonance Theory (Festinger, 1957), users experiencing a mismatch between their expectations and observed system behavior may adjust their actions to reduce discomfort, altering their decision-making and interaction patterns. To explore this further, we hypothesize: H6: When users believe that a model has switched, it will lead to changes in their behavior.
@@ 3.2. Experiment Setup
To explore these hypotheses, we designed an online survey experiment where participants interacted with a facial recognition-based web application (see Figure 1), modeled after the search results page on CWPS (Mohanty et al., 2019b). In each trial, participants were shown a query photo and tasked with identifying potential matches (i.e., other photos of the same person) from a ranked list of search results retrieved by one of two distinct facial recognition models (see Section 3.3 for details). The assignment of models was randomized behind the scenes, and participants were not informed which model was being used or if the model changed between trials. Participants visually inspected the results to determine whether any of the retrieved images matched the person in the query photo.
@@
Each participant completed 8 trials, with each trial featuring a different query photo and its corresponding search results (i.e., potential candidates). This randomization ensured that, while participants were unaware of the underlying model, we could systematically explore their ability to detect changes between the models and the cues they relied on. This design enabled us to examine shifts in user perceptions, performance, and behaviors across trials.
3.3. Facial Recognition Models
Similar to CWPS (Mohanty et al., 2019b), we used facial recognition models from Microsoft Azure for this study: the oldest available model, recognition_01 (released in 2017), and the most recent model, recognition_04 (released in 2021). For clarity, we will refer to these as the old model and new model, respectively throughout this paper. According to the developer website (Urban, 2023), the new model is supposed to be the most accurate available, designed to handle challenging cases like faces with masks or facial hair. However, its effectiveness in specialized domains, such as historical photo identification, remains unclear. Previous benchmarking of the old model on CWPS showed that while effective at retrieving correct matches, it also retrieved a large number of search results (mostly false positives), increasing the likelihood of incorrect matches (Mohanty et al., 2020).
We set the confidence threshold for retrieving results at 0.50 based on prior work (Mohanty et al., 2020), which demonstrated its utility in avoiding false negatives (i.e., missing the correct match) while ensuring a manageable set of retrieved results for the older model. However, given the newer model’s substantially improved precision and recall (Appendix A.2), this threshold produced significantly fewer results: an average of 20 compared to 378 with the older model. While this large discrepancy in result count might suggest a system-level design difference, it actually reflects the inherent capabilities of the models under identical configurations. To preserve these native configurations and better simulate real-world usage scenarios, we conducted our study without artificially constraining result counts. However, we acknowledge this as a potential limitation in Section 6.
Our benchmarking study (see details in Appendix  A.2) on the dataset we used for the study (see Section 3.5) revealed notable differences between the two models: the new model consistently outperformed the old model in terms of average precision (i.e., ranking correct matches higher); 0.79 (new) vs. 0.41 (old) and recall (i.e., finding more correct matches); 0.76 (new) vs. 0.54 (old). Additionally, while the old model retrieved a large number of results, many were false positives, making it relatively less precise in this domain. In contrast, the new model produced fewer results overall but retrieved more correct matches with higher confidence scores.
These differences in key metric s— such as the number of results, the ranking of more similar-looking candidates, and improved handling of diversity — make the old model and new model compelling candidates for this study. However, while these measurable differences exist, an important question remains: can users actually perceive these differences in practice? Since participants were unaware of which model was in use during each trial, the study explores whether users can rely on visible cues, such as fewer but more accurate results and better-ranked matches, to distinguish between the two models.
3.3.1. Latency in retrieving results
Prior work has demonstrated that latency, i.e., the time taken by a system to retrieve result, can significantly shape user perceptions of algorithmic systems, where even slight delays reduce engagement and satisfaction in search and chatbot contexts, but in some cases, slower responses may foster reflection or higher perceived quality (Park et al., 2019; Teevan et al., 2013; Brutlag, 2009; Gnewuch et al., 2018). In this study, we introduced an artificial delay to explore whether users consider latency when trying to distinguish between model changes and how it may influence their perception of model accuracy. Participants were assigned to one of four conditions: (i) Old Model retrieves results faster than New Model, (ii) New Model retrieves results faster than Old Model, (iii) both models retrieve results with random latency, or (iv) both models retrieve results with identical latency. For slower conditions, we introduced a 12-second delay, while standard conditions took 3 to 5 seconds.
3.4. Interface
We designed the web application to guide participants through the person identification tasks across eight consecutive trials. In each trial, participants were shown a ranked list of search results for the query photo, retrieved by one of the two models described in Section 3.3 (the old model or the new model). The specific model for each trial was randomly assigned, and participants were blinded to the model’s identity (no old/new labels were shown) or whether the model had changed between trials. This ensured that participants’ assessment of the models were based solely on their interaction with the model outputs, without any influence from model labels or associations. The interface supported five key components:
  1. (1)